The specification on a heater coil is 250 V, 500 W. Calculate the resistance of the coil. What will be the resistance of a coil of 1000 W to operate at the same voltage?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. (a) The power consumed by a coil of resistance R when connected across a supply v is P = v^2/R The resistance of the heater coil is, therefore R = v^2/P = (250)^2/500 = 125 Ω (b) If P = 1000 w then R = v^2/P = (250)^2/1000 = 62.5 Ω